Tom exerts a 250 N force on the shot at an angle of 75° above horizontal. The mass of the shot is 7 kg. How big is the resultant force? What is the angle of this resultant force with respect to horizontal?
Answer #1

Gravitational force on shot = mg = 7*9.8 = 68.8 N

Angle between two forces = 90+75 = 165 degree

Resultant = sqrt(250^2+68.8^2+2*250*68.8*cos 165 degree)

= 184.4 N answer

Angle with horizontal = arctan((250*sin 75 degree - 68.8)/(250 cos 75 degree))

= 69.46 degree above horizontal